A rocket of mass 20 kg has 180 kg of fuel.The exhaust velocity of fuel is 1.6km/sec.Calculate the ultimate velocity of the rocket gained,when the rate of consumption of fuel is 2kg/sec.

Grade:11

3 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear ramya bhaskar

Velocity of rocket is given as

v=u ln(M/M-rt)

wher M is the total mass of rocket + fuel at t=0 sec = 20 +180 =200 kg

and r rate of fuel consumption =2kg/sec

u is the velocity of gas ejection w.r.t rocket = 1.6 km/sec

 

ultimate velocity is achieved when all the fuel will burn

ie. M-rt =20 kg

so v= 1.6 ln(180/20)

        =3.515 km/sec


Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin



Pavithra
5 years ago
A rocket having initial mass of 240kg ejects fuel at the rate of 6kg/s with a velocity of 2km/s vertically downward relative to itself. Calculate its velocity 25seconds after start taking initial velocity to be zero and neglecting gravity
 
KUMARI
5 years ago
Sjxjwjdjxjsudusududuchshcushchshchshcuejfkckdkcjdjckdncjsjdjsjcjsjcjririeufue8t7f7eu27r8fufjeirfivjrieigicjeir8viejvidieiejc8eue8t8ci2ir848t8f82747f7wudovo29rodcjiefjwjcjejfjcj